OpenStudy (anonymous):

How do I simplify √50/9 ?

OpenStudy (anonymous):

\[\sqrt{50} = \sqrt{25*2} = \sqrt{25}*\sqrt{2} = 5\sqrt{2}\]

OpenStudy (mathstudent55):

\( \sqrt{ \dfrac{50}{9} } \) First, separate the roots: \( \dfrac{ \sqrt{50} } { \sqrt{9}} \) Now factor the largest perfect-square factor out of 50. Follow what @Euler271 did above for the numerator. The denominator is simply sqrt(9).